OpenStudy (anonymous):

prove that \[(ab)^{-1} = a^{-1}b^{-1}\] prove that \[(ab)^{-1} = a^{-1}b^{-1}\] @Mathematics

myininaya (myininaya):

\[(ab)^{-1}(ab)=1\] \[aa^{-1}b^{-1}b=1\] then this means \[(ab)^{-1}(ab)=aa^{-1}b^{-1}b\] multiplication is commutative so we can write \[(ab)^{-1}(ab)=a^{-1}b^{-1}ab\] \[(ab)^{-1}(ab)=a^{-1}b^{-1}(ab)\] since ab=ab then this must mean that \[(ab)^{-1}=a^{-1}b^{-1}\]

OpenStudy (anonymous):

my book game me this one-liner:\[ab(a^{-1}b^{-1}) = (a*a^{-1})(b*b^{-1}) = 1; \ \ hence\ \ a^{-1}b^{-1} = (ab)^{-1}\]

myininaya (myininaya):

looks similar to mine

OpenStudy (anonymous):

but he started a little different.

myininaya (myininaya):

they did the whole equals to one thing

OpenStudy (anonymous):

\[1 = ab*(ab)^{-1}\]